How many grams of potassium carbonate does 1.6L of a .29M solution contain?
1
Expert's answer
2019-05-22T07:03:42-0400

c(K2CO3) = 0.29M = 0.29 mol/dm3

V(K2CO3) = 1.6 dm3

n(K2CO3) = cV = 1.6 * 0.29 = 0.464 mol

m(K2CO3) = n/M = 0.464*138 = 64.032 g